OpenStudy (anonymous):

Find the equation of the locus of a point p which moves such that the sum of the squares of its distances from (3,0) and (-3,0) is 68

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

hi!!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i guess we can do this, it is gong to take a bit of algebra

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

lets call such a point \((x,y)\) and make a equation

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

the square of the distance between \((x,y)\) and \((3,0)\) is \((x-3)^2+y^2\)

OpenStudy (misty1212):

the square of the distance between \((x,y)\) and \((-3,0)\) is \[(x+3)^2+y^2\]

OpenStudy (misty1212):

add them together, set the result equal to \(68\) get \[(x-3)^2+(x+3)^2+2y^2=68\] now more algebra

OpenStudy (anonymous):

ok so x^2-6x+9+y^2 and x^2+6x+9+y^2 so adding them together we get 2x^2+18+2y^2 like that?

OpenStudy (misty1212):

that is what i get, yes

OpenStudy (misty1212):

oh equal 68

OpenStudy (misty1212):

\[2x^2+18+2y^2 =68\] subtract \(18\) maybe to get \[2x^2+2y^2=50\] then divide by \(2\) to make it look neater

OpenStudy (misty1212):

\[x^2+y^2=25\] might might be a nicer answer a circle with center \((0,0)\) and radius \(5\)

OpenStudy (anonymous):

so thats the equation of the locus

OpenStudy (anonymous):

??

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

assuming we did all the algebra right looks good to me

OpenStudy (anonymous):

but we didnt show that thee distance is 68..how would we do that?

OpenStudy (misty1212):

the sum of the square of the distances was supposed to be 68 right?

OpenStudy (anonymous):

yea

OpenStudy (misty1212):

\[(x-3)^2+(x+3)^2+2y^2\] was the sum of the square of the distances \[(x-3)^2+(x+3)^2+2y^2=68\] says it is equal to 68

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

then the algebra turned that in to \[x^2+y^2=25\]

OpenStudy (anonymous):

ok i understand now thank you

OpenStudy (anonymous):

but ...what happen to the square root sign

OpenStudy (anonymous):

\[(\sqrt{(x-3)^2+(y-0)^2})^2+(\sqrt{(x+3)^2+(y-0)^2})^2=68\]ok i see now

OpenStudy (anonymous):

yea thank you !!

OpenStudy (misty1212):

lol \[\color\magenta\heartsuit\]